OpenStudy (anonymous):

How do I find X by substitution? 2x + y = 1 4x + 2y = −2

OpenStudy (dan815):

2x + y = 1 <---- rearrange for y=.... and plug that ..... into ---- V 4x + 2y = −2

OpenStudy (anonymous):

after I change it into 4x + 2(2x +1) -2 4x + 4x + 3 = -2

OpenStudy (anonymous):

do i combine like terms then simplify?

OpenStudy (dan815):

u doing something wrong

OpenStudy (dan815):

remember y = 1-2x

OpenStudy (dan815):

u shud get 2 = -2 which cannot be possible, so there are no solutions to this equation

OpenStudy (dan815):